OpenStudy (anonymous):

2. How can you tell when a quadratic equation has two identical, rational solutions? (1 point) when the radicand is negative when b in the quadratic formula is greater than the radicand when the radicand equals zero when the radicand is not a perfect square

OpenStudy (anonymous):

the second one. When B in the formula is greater. No perfect square quadratics can have be smaller than the radicand. I think.

OpenStudy (albert0898):

When b^2=4ac the zero is at a single point. The parabola does not penetrate the x axis. This represents two zeros, both with the same value. Example: x^2+4x+4 yields two zeros, each at x=-2. Notice that b^2=16=4ac Penetrate = succeed in forcing a way into or through Parabola = a symmetrical open plane curve formed by the intersection of a cone with a plane parallel to its side. The path of a projectile under the influence of gravity ideally follows a curve of this shape.

OpenStudy (anonymous):

so it is b?

OpenStudy (albert0898):

Mmmm... I don't know. I would just BUMP this and see if people say B or other answers. Just post this question in the chat.

OpenStudy (anonymous):

it has to be b

OpenStudy (mertsj):

When the radicand is 0 because, watch this:

OpenStudy (mertsj):

\[x=\frac{-b \pm \sqrt{0}}{2a}\]

OpenStudy (albert0898):

I think it's C.

OpenStudy (mertsj):

Would you agree that -b+0 is the same as -b+0?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So if the radicand is 0, there will only be one answer.

OpenStudy (mos1635):

quadratic equation: ax^2+Bx+c=0 is that it???

OpenStudy (mertsj):

yes

OpenStudy (mos1635):

identical, rational solutions: X1=X2 both \[\in R\] ???

OpenStudy (mertsj):

That means the radicand is 0 because the "two roots" are equal.

OpenStudy (mos1635):

C it is